OpenStudy (anonymous):

Solve y over negative 6 + 5 = 9

OpenStudy (tanya123):

The way you solve an equation like this is by pretending y is a number. If y were a number and you had to calculate the left side, you would 1. first divide y by -6 2. then you would add 5. You must use the order of operations. In order to solve the equation, you need to undo what is being done to y. To undo it, you do the opposite operation in the opposite order. To do the opposite order, we start with step 2 then we do step 1. Start by step 2 above and undo it. That means instead of adding 5, you subtract 5. Remember, you must do the same to both sides of the equation, so subtract 5 from both sides. y−6+5−5=9−5 y−6=4 Now we do step 1 in reverse. Step 1 is divide y by -6, so multiply both sides by -6. y−6×(−6)=4×(−6) y=−24 The solution is: y = -24

OpenStudy (the_fizicx99):

Is it \(\ \sf \dfrac{y}{-6 + 5} = 9 \) or \(\ \sf \dfrac{y}{6} + 5 = 9\) ?